OpenStudy (anonymous):

Which of the following has a graph that is a straight line? Equation 1: y = 4x3 + 6 Equation 2: y = 5x - 4.5 Equation 3: y2 = x - 1 Equation 4: y = 2x2 + 6

OpenStudy (anonymous):

All except eq.3

OpenStudy (anonymous):

1 and 4 do not assuming that 4x3 = 4x^3 and 2x2 = 2x^2

OpenStudy (anonymous):

so would it be 2?

OpenStudy (anonymous):

both 2 and 3 make straight lines, but have different slopes. Can you double check those to make sure you typed them correctly?

OpenStudy (anonymous):

I copy and pasted them

OpenStudy (anonymous):

okay wait I missed the y^2 that makes a circle I believe so yes 2 would be correct .

OpenStudy (anonymous):

1 is supposed to be Y= 4X^3+6

OpenStudy (anonymous):

2 is Y=5X-4.5

OpenStudy (anonymous):

3 is y^2=x-1

OpenStudy (anonymous):

and 4 is y=2x^2+6

OpenStudy (anonymous):

1 is a cubic so not a straight line 2 is a straight line 3 is a circle 4 is a parabola

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Can you help mewith two more?

OpenStudy (anonymous):

me with*

OpenStudy (anonymous):

I missed the Y^2 in 3 originally. No problem! =)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which statement best explains whether the table represents a linear or nonlinear function? Input Output 0 1 1 2 2 4 3 16 It is a linear function because there is a constant rate of change in both the input and output values. It is a nonlinear function because there is a constant rate of change is both the input and output values. It is a linear function because the output values are increasing at different rates. It is a nonlinear function because the output values are increasing at different rates.

OpenStudy (anonymous):

hmm it is definitely linear because the values only increase. I am thinking it is th first one

OpenStudy (anonymous):

One second let me develop an equation for this table

OpenStudy (anonymous):

I was wrong it is not linear because the slope is changing.

OpenStudy (anonymous):

so it would be the fourth one

OpenStudy (anonymous):

That table would make a quadratic I believe.

OpenStudy (anonymous):

Ok last one...

OpenStudy (anonymous):

Which of the following ordered pairs could be placed in the table below and still have the relation qualify as a linear function? Input Output −5 7 −3 15 −1 23 ? ? (0, 31) (−7, 30) (1, 31) (0, −1)

OpenStudy (anonymous):

Since the slope changes between the points it means that there must be and x^n in the equation and n cant = 1 which means it is not linear. Slope = rate of change.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

so the points are: (5,7) (3,15) (1,23) So we need to develop an equation for these points 7-15 -8 ---- = ---- = -4 5-3 2 So the slope is -4 y-7=-4(x-5) y-7=-4x+20 y=-4x+27 using that equation however none of the points given work they are: (0,31) (7,30) (1,31) (0,1) correct?

OpenStudy (anonymous):

@wabbitshark those are wrong! I just used them and I got a D on my test LOL :(